Circular No. 5870

Central Bureau for Astronomical Telegrams INTERNATIONAL ASTRONOMICAL UNION

Postal Address: Central Bureau for Astronomical Telegrams Smithsonian Astrophysical Observatory, Cambridge, MA 02138, U.S.A. Telephone 617-495-7244/7440/7444 (for emergency use only) TWX 710-320-6842 ASTROGRAM CAM EASYLINK 62794505 MARSDEN@CFA or GREEN@CFA (.SPAN, .BITNET or .HARVARD.EDU)

NOVA LUPI 1993 M. Della Valle, European Southern Observatory, reports: “A series of

spectrograms (range 390-1000 nm, resolution 1-2 nm) of N Lup 1993 were obtained on Sept. 23.96 UT with the New Technology Telescope (+ EMMI) at La Silla. The spectrum shows emission lines of H, Fe II, Na I, O I, and N I, indicating that this nova has been caught during its early decline. The expansion velocity derived from the FWHM of Hα is ∼ 2700 km/s.”

SUPERNOVA 1993Y IN UGC 2771 J. Mueller reports her discovery of an apparent supernova of mag ∼ 17,

located 9′′ east and 37′′.6 north of the nucleus of UGC 2771 (α = 3^{h}28^{m}.1, δ = +39^{o}35′, equinox 1950.0). The IIIa-J discovery plate was taken by Mueller on Sept. 18 UT with the 1.2-m Oschin Telescope in the course of the second Palomar Sky Survey.

A spectrogram obtained on Sept. 22 UT by A. Dressler with the 5m Hale Telescope (+ 4-shooter + spectrograph), analyzed by W. Sargent, shows this to be a type-Ia supernova, 20-21 days past maximum. A. V. Filippenko and T. Matheson, University of California at Berkeley, also report that CCD spectra (range 310-1000 nm) obtained on Sept. 25 with the Shane 3-m reflector at Lick Observatory show that the supernova is of type Ia, but closer to 1 month past maximum brightness.

SUPERNOVA 1993Z IN NGC 2775 R. R. Treffers, A. V. Filippenko, B. Leibundgut, Y. Paik, and L. F. M.

Lee, University of California at Berkeley; and M. W. Richmond, Princeton University, report their discovery of a supernova in NGC 2775 (α = 9^{h}07^{m}41^{s}, δ = +7^{o}14′.5, equinox 1950.0), ∼ 15′′ west and 42′′ south of the galaxy’s nucleus. The object (magnitude R = 13.9 ± 0.3) was found on Sept. 23 UT during the Leuschner Observatory Supernova Search, which uses an automated 0.76-m telescope equipped with the Lawrence Berkeley Laboratory CCD camera. Confirmation was obtained on Sept. 24 at about the same brightness. It could not be detected to a limiting magnitude of R = 18.3 in images obtained on May 22. Spectra (range 310-990 nm, resolution 0.6 nm) obtained on Sept. 25 with the Shane 3-m reflector at Lick Observatory reveal that the supernova is of type Ia, ∼ 4 weeks past maximum brightness.

1993 September 25 Daniel W. E. Green